# Operational research approaches and mathematical models for kidney exchange: A literature survey and empirical evaluation

### v1

@@ 2.4 Analyses Based on Dynamic KEPs
We now consider the literature on methods that attempt to maximise the number of transplants over a time period, as it may not coincide with maximising the number of transplants at each matching run (also known as the myopic approach). To avoid the underutilisation of RDPs in a current matching run, Dickerson et al., 2012a used the composition of previous matching runs (using generated datasets) to learn the potential of each node, which is defined as the expected contribution of that node to the objective function. They then determine, at each matching run, a set of exchanges that maximises the number of transplants minus the potential of the nodes included in the solution. High-potential nodes (such as NDDs with blood group O) are thus typically included in the solution only when they enable multiple transplants involving low-potential nodes, avoiding their underutilisation. Their simulations show that learning a set of only 20 node potentials (one for each blood group combination of RDPs and NDDs) is already sufficient to significantly increase the total number of transplants compared to the myopic approach while also being computationally tractable. Carvalho et al., (2024) adopted a similar strategy in the context of the Canadian Kidney Paired Donation Program. They used a set of 80 node potentials (this time based on blood groups and tissue-type compatibility) and showed that the resulting method produced solutions that were better compared to the myopic approach in terms of the number of transplants, average waiting time, and their measure of fairness. Awasthi and Sandholm, (2009) used a related strategy in which they attributed a score to each exchange, which was computed through a scenario-based algorithm. Their results also suggest that a scoring method outperforms the myopic approach, even though it can be computationally challenging to apply such a method to larger instances, given the resulting large number of possible exchanges (and hence, scores to approximate). Chisca et al., 2019b built upon the work of Awasthi and Sandholm, (2009) with the objective of improving scalability.
Anshelevich et al., (2013) discussed a weighted dynamic KEP and observed that the myopic approach often included transplants with low weights, which could be seen as an underutilisation of the nodes involved in such transplants. They suggested that adopting a threshold strategy – based on hiding all transplants with weight below a certain threshold – could be highly beneficial.
@@
One particularly relevant application of online matching problems is the consideration of deceased-donor initiated chains (DDIC). In many countries, chains triggered by NDDs are terminated by the final donor of the chain donating to a recipient on the DDWL. In Italy, however, chains can also be initiated by deceased donors – leading to DDICs (Furian et al., , 2019, 2020). One major consideration when dealing with DDICs is the fact that the first transplant of the chain cannot wait for the next matching run. Whereas Cornelio et al., (2019) argued that DDICs could benefit both the recipients on the DDWL and the recipients within the KEP pool, Wall et al., (2017) highlighted that DDICs also raise several ethical issues.
@@
When considering dynamic models with periodic matching runs, one natural question is how frequently should matching runs take place? Whereas triggering a matching run each time an RDP or an NDD joins the pool is a possibility, waiting instead for more RDPs and NDDs to arrive to thicken the pool (Roth, , 2008) is another. One could think that the latter is always more advantageous than the former as it naturally leads to more transplants, but this does not account for the departure rate: some RDPs and NDDs leave the pool without being matched, for example, due to a deterioration of their health condition. In addition, the total number of transplants is not the only metric that is relevant in KEPs; other important metrics include the average waiting time.
@@
Ashlagi et al., (2018) ran simulations on the data of two KEPs in the USA, this time considering a departure rate. They showed that matching frequently does not harm the fraction of transplanted RDPs, whereas matching infrequently may result in the departure of easy-to-match RDPs. These conclusions were supported by the experiments of Ashlagi et al., (2023), who studied the correlation between thicker pools and the number of transplants in the presence of two RDP groups: hard-to-match and easy-to-match. Therefore, one reason for not waiting for the pool to thicken is the potential loss of RDPs who are close to leaving the pool. Akbarpour et al., (2020) studied the case in which one could accurately predict RDPs’ departure times and demonstrated that significant gains could be achieved if such a prediction were available. In the absence of accurate predictions, however, they also concluded that online matching is preferable over waiting for the pool to thicken.
@@
Allowing longer chains typically leads to better outcomes for KEPs, producing exchanges with more transplants that also benefit highly sensitised patients (Ashlagi et al., , 2012; Anderson et al., , 2015). Longer chains are possible from a logistical point of view because the transplants involved in a chain can be performed non-simultaneously. In the dynamic setting, relaxing the simultaneity constraint enables non-simultaneous extended altruistic donor (NEAD) chains, in which the final donor in a chain segment becomes an NDD in the following matching run. Such an NDD is sometimes referred to as a bridge donor. Some early NEAD chains were reported by Rees et al., (2009). One of the NEAD chains that was reported consisted of ten kidney transplantations coordinated over a period of eight months initiated by a single NDD.
@@
In theory, NEAD chains are beneficial (Ashlagi et al., , 2013). In practice, however, one must also account for the possibility that a bridge donor reneges (or departs), which becomes more likely as the time between two matching runs increases. Note that reneging cannot normally be prevented by legal means as, in most countries, it is not possible to make organ donation a legally binding obligation (Dickerson et al., 2012a, ). Both Gentry et al., (2009) and Ashlagi et al., 2011b studied dynamic KEPs considering a reneging rate for the bridge donors and obtained conflicting results. The former concluded that allowing NEAD chains decreased the overall number of transplants due to reneging risks. In contrast, the latter found that this was not the case and that, in fact, NEAD chains increased the number of transplants, including for highly sensitised recipients. The two sets of authors discussed the matter further (Ashlagi et al., 2011a, ; Gentry and Segev, , 2011), suggesting that differences in experimental setups between the two studies were likely the cause of their divergent conclusions. This is a good time to remind the reader that many of the conclusions drawn from empirical experiments in the KEP literature are, indeed, highly dependent on the modelling assumptions used in those experiments, a fact that is usually acknowledged by the authors of such studies.
2.5 Robustness in KEPs
@@
Thus far, we have focused on how to select a set of exchanges for transplantation, yet in practice, some identified transplants will not proceed. In the UK’s national KEP from 2019 to 2023, around $69\%$ of selected transplants proceeded to surgery (NHS Blood and Transplant, , 2023). There are many reasons why a selected exchange may not proceed, including vertex failure (e.g., a donor or recipient becoming ill) or arc failure (e.g., a laboratory crossmatch being identified). This section gives an overview of research into robust optimisation in KEPs. The first direction studies stochastic approaches based on the expected likelihood of a transplant proceeding, whilst the second line of research is concerned with recourse, focusing on methods to recover transplants when parts of exchanges can no longer proceed.
@@
One proposal to account for the inevitable loss of transplants between their selection and their realisation is to consider their likelihood of failure when selecting them. A possible way of doing this is to maximise the expected number of transplants in a stochastic setting, as initially suggested by Dickerson et al., (2013) and Pedroso, (2014). The model given by Dickerson et al., (2019) provides each node and arc with a probability of success, from which a set of likely exchanges can be selected. Their branch-and-price algorithm identifies a solution that has maximum expected utility. Utilisation of such an objective can raise issues with highly-sensitised recipients not being selected, hence, Dickerson et al., (2019) extended their study to a dynamic setting that yields better outcomes for such recipients. They also considered the use of partially successful paths within a selected chain. Namely, when there is some failure in a chain, then the initial part of the chain can proceed. Goldberg and Poss, (2022) extended upon this, giving a mixed-integer linear programming model for partially successful paths within a selected chain.
A different branch-and-price algorithm was provided by Alvelos et al., (2019), however, in the case when the probability of failure is equal among all nodes and equal among all arcs. Bidkhori et al., (2020) also extended the work of Dickerson et al., (2019), providing a mixed-integer linear programming reformulation that is compact (i.e., the numbers of variables and constraints are polynomial in the input size) when the probability of failure is inhomogeneous among arcs. This is unlike the model of Dickerson et al., (2019), which requires the enumeration of all feasible cycles and chains, which can be intractable even for small exchanges. Zheng et al., (2015) took a different approach by modelling KEPs as a stochastic minimum cost flow problem.
@@
Smeulders et al., 2022a studied the use of laboratory crossmatch tests on some potential transplants before a matching run is performed, to give certainty that those transplants may proceed. Their model thus comprises two steps. The first step selects a set of potential transplants for laboratory crossmatching. With this certainty from the laboratory tests, they then found the set of exchanges that maximised the expected number of transplants performed.
@@
Although internal recourse has been shown to improve the number of transplants that proceed in practice, other methods of recourse have been developed. For example, Klimentova et al., (2016) considered the notion of subset-recourse, where vertices for a recourse exchange can involve vertices remaining from a failed exchange as well as vertices not initially selected for transplant. Carvalho et al., (2021) studied different recourse policies and provided integer programming models for each policy; their full recourse model finds a set of exchanges that maximises the number of the originally selected RDPs. Blom et al., 2024a studied recourse in terms of a three-stage defender-attacker-defender model, where the KEP organisers (the defender) select a set of exchanges, then the most disruptive set of RDPs and NDDs withdraw after being selected in a matching run (replicating an attacker or informed adversary), which is followed by the KEP repairing the solution using only the remaining RDPs and NDDs. They used a cutting plane method for the latter two stages of the problem, which allowed their observed running times to outperform the model from Carvalho et al., (2021). Chisca et al., 2019a took a more combinatorial approach that assessed if an alternative solution could be found with minimal changes to the original selection.
@@ 2.6 Multi-Hospital and International KEPs
Klimentova et al., (2021) introduced a different credit system based on the assumption that the agents are not strategic. At a given “round” (matching run), each country has a target number of kidney transplants, representing a “fair” allocation. The difference between the actual number of transplants for a country and its target number is then used to update that country’s number of credits (positively or negatively), and the credit balances are then used to adjust the target allocations for the next round. Biró et al., (2020) conducted simulations in relation to this credit-based framework, whilst Benedek et al., 2024b extended these simulations to a larger number of countries (for cycles of length $2$ only), and for a range of different ways of computing the initial set of exchanges. At each subsequent round, they compute a solution with the maximum number of transplants that lexicographically minimises the deviations from the target allocations for each country. The credit-based framework was extended to the case of unbounded length exchanges by Benedek et al., 2024a . Associated complexity results relating to these credit-based frameworks can be found in Biró et al., 2019b and Benedek et al., (2023, 2025).
@@
When the agents represent countries collaborating in an international KEP, additional logistical constraints may be required, such as countries permitting different lengths of cycles and chains due to differences in legislation between the countries. Mincu et al., (2021) presented an integer programming model that handles these constraints in the context of an international KEP.
@@ 2.7 Recipients’ Preferences
In a KEP, the suitability of a donor for a given recipient is generally modelled by the presence of an arc in the underlying compatibility graph and its associated cardinal weight. Indicators of the utility of a potential transplant that can contribute to this weight can include, for example, the level of HLA-matching between donor and recipient, the age of the donor and the waiting time of the recipient. Instead of cardinal utilities, an alternative approach is to allow recipients to express ordinal preferences over their potential donors. In the presence of ordinal preferences, the aim is typically to find a stable set of exchanges $S$ (comprising cycles and chains), meaning that there is no blocking exchange, i.e., an exchange $E$ such that each recipient in $E$ prefers their donor in $E$ to the donor they receive in $S$ (if any).
@@
Biró and McDermid, (2010) defined the $b$-way stable $l$-way exchange problem, which is the variant of KE-Opt with ordinal recipient preferences where we seek a set of exchanges comprising cycles of length at most $l$, such that there is no blocking cycle of length at most $b$. In the case that $b=l=2$, we obtain the classical Stable Roommates problem (Irving, , 1985), as observed by Roth et al., (2005). Hence, Irving’s algorithm (Irving, , 1985) can be used to find a stable set of exchanges or report that none exists in linear time. If $b=l=\infty$, we obtain the classical Housing Market problem as observed by Roth et al., (2004), for which a stable set of exchanges always exists and can be found in linear time using Gale’s Top Trading Cycles Mechanism (Shapley and Scarf, , 1974). On the other hand, when $b=l=3$Biró and McDermid, (2010) showed that the problem of deciding whether a stable set of exchanges exists is $\mathcal{NP}$-complete. Irving, (2007) showed that $\mathcal{NP}$-completeness also holds in the case that $b=3$ and $l=2$, whilst Mészáros-Karkus, (2017) proved an analogous result for the case that $b=2$ and $l=3$. Moreover, the author also showed that the problem is $\mathcal{W}[1]$-hard when parameterised by the number of cycles of length $3$.
In the same setting as Biró and McDermid, (2010), Huang, (2010) independently studied three notions of stability in KEPs that differ in their restrictiveness, namely weak, strong, and super stability. (Weak stability corresponds to stability as defined informally above, and strong stability was also introduced by Biró and McDermid, (2010)). Huang, (2010) showed that the $3$-way stable $3$-way exchange problem is $\mathcal{NP}$-complete for each of these notions of stability. Moreover, for so-called strong stability, Huang, (2010) showed that counting the number of sets of exchanges satisfying this property is a #$\#\mathcal{P}$-complete problem.
Klimentova et al., (2023) gave four ILP models for the problem of finding a stable set of exchanges or reporting that none exists. They conducted simulations, measuring computation times and how many instances did not admit a stable set of exchanges. They also explored the trade-off between size (number of transplants in a set of exchanges) and stability (regarding the number of blocking exchanges). Baratto et al., (2025) defined a new stability concept, which they called local stability, where a blocking exchange must contain at least one vertex from the initial set of selected exchanges. They proposed an ILP model to find a locally stable set of exchanges, and via simulations, they found that non-empty locally stable sets of exchanges frequently exist, even in compatibility graphs that do not admit a stable set of exchanges.
@@
Cardinal preferences are also used in KEPs: they are most commonly represented by assigning weights to potential transplants that reflect their utility. Freedman et al., (2020) created a tie-breaking scoring function determined by a KEP’s stakeholders. Their opinions were elicited via a series of pairwise comparisons between recipient profiles, which determines what should be prioritised. Their preliminary testing of whether stakeholders’ opinions could be incorporated showed that RDPs with underdemanded blood group combinations would be negatively impacted the most. In contrast, many other RDPs’ chances would remain unchanged. The process of querying stakeholders was also taken by McElfresh et al., (2020), who proposed querying certain donors and recipients to check if they would accept a particular transplant before the selection phase. They intended to minimise rejections of selected transplants post-selection (relating to Section 2.5 on robustness in KEPs). Dickerson and Sandholm, (2015) proposed a similar model that learns high-level objectives from experts, and their framework implements them based on previous matching run data.

### v2

@@ 2.4 Analyses Based on Dynamic KEPs
We now consider the literature on methods that attempt to maximise the number of transplants over a time period, which need not coincide with maximising the number of transplants at each matching run (also known as the myopic approach). The approach of Awasthi and Sandholm, (2009) attributed a score to each exchange, which reflects the expected weight (or utility) of that exchange and was computed through a scenario-based algorithm. Their results suggest that a scoring method outperforms the myopic approach, even though applying such a method to larger instances can be computationally challenging, given the resulting large number of possible exchanges (and hence, scores to approximate). Chisca et al., 2019b built upon the work of Awasthi and Sandholm, (2009) with the objective of improving scalability.
In related work that aims to avoid the underutilisation of RDPs in a current matching run, Dickerson et al., 2012a used the composition of previous matching runs (using generated datasets) to learn the potential of each node, i.e., the expected contribution of that node in the solution. At each matching run, they determine a set of exchanges that maximises the number of transplants minus the potential of the nodes included in the solution. High-potential nodes (such as NDDs with blood group O) are typically selected only when they enable multiple transplants involving low-potential nodes, avoiding their underutilisation. Their simulations demonstrate that their approach can significantly increase the total number of transplants compared to the myopic approach, while also being computationally tractable. Carvalho et al., (2024) gained similar results when they adopted an analogous strategy in the context of the Canadian Kidney Paired Donation Program, showing improvements in terms of the number of transplants, average waiting time, and their measure of fairness.
Anshelevich et al., (2013) discussed a weighted dynamic KEP and observed that the myopic approach often included transplants with low weights, which could be seen as an underutilisation of the nodes involved in such transplants. They suggested that adopting a threshold strategy – based on hiding all transplants with weight below a certain threshold – could be highly beneficial. Finally, we note that Ashlagi et al., (2019) analysed more general dynamic matching markets, especially in relation to hard-to-match agents, with reference to kidney exchange.
@@
One particularly relevant application of online matching problems is the consideration of deceased-donor initiated chains (DDIC). In many countries, chains triggered by NDDs are terminated by the final donor of the chain donating to a recipient on the DDWL. In Italy, however, chains can also be initiated by deceased donors – leading to DDICs (Furian et al., , 2019, 2020). One major consideration when dealing with DDICs is the fact that the first transplant of the chain cannot wait for the next matching run. Cornelio et al., (2019) and Billa et al., (2018) argued that DDICs could benefit both recipients on the DDWL and within the KEP pool, and Wang et al., (2022) showed via simulations that using only 3% of deceased donors in DDICs could have a large positive impact on the KEP. However, Wall et al., (2017) highlighted that DDICs also raise several ethical issues.
@@
When considering dynamic models with periodic matching runs, one natural question is how frequently should matching runs take place? Whereas triggering a matching run each time an RDP or an NDD joins the pool is a possibility, waiting instead for more RDPs and NDDs to arrive to thicken the pool (Roth, , 2008) is another. One could think that the latter is always more advantageous than the former as it naturally leads to more potential transplant opportunities, but this does not account for the departure rate: some RDPs and NDDs leave the pool without being matched, for example, due to a deterioration of their health condition. In addition, the total number of transplants is not the only metric that is relevant in KEPs; other important metrics include the average waiting time.
@@
Whilst the papers surveyed in the previous paragraph do not consider RDPs leaving the pool without a transplant, the following papers do take departure rates into account, which can lead to different conclusions. Ashlagi et al., (2018) ran simulations on the data of two KEPs in the USA, when considering a departure rate. They showed that matching frequently does not harm the fraction of transplanted RDPs, whereas matching infrequently may result in the departure of easy-to-match RDPs. These conclusions were supported by the experiments of Ashlagi et al., (2023), who studied the correlation between thicker pools and the number of transplants in the presence of two RDP groups: hard-to-match and easy-to-match. Therefore, one reason for not waiting for the pool to thicken is the potential loss of RDPs who are close to leaving the pool. Akbarpour et al., (2020) studied the case in which one could accurately predict RDPs’ departure times and demonstrated that significant gains could be achieved if such a prediction were available. In the absence of accurate predictions, however, they also concluded that online matching is preferable over waiting for the pool to thicken.
@@
Allowing longer chains typically leads to better outcomes for KEPs, producing exchanges with more transplants that also benefit highly sensitised patients (Ashlagi et al., , 2012; Anderson et al., 2015b, ). Longer chains are possible from a logistical point of view because the transplants involved in a chain can be performed non-simultaneously. In the dynamic setting, relaxing the simultaneity constraint enables non-simultaneous extended altruistic donor (NEAD) chains, in which the final donor in a chain segment becomes an NDD in the following matching run. Such an NDD is sometimes referred to as a bridge donor. Some early NEAD chains were reported by Rees et al., (2009). One of the NEAD chains that was reported consisted of ten kidney transplantations coordinated over a period of eight months initiated by a single NDD.
@@
In theory, NEAD chains are beneficial (Ashlagi et al., , 2013). In practice, however, one must also account for the possibility that a bridge donor reneges (or departs), which becomes more likely as the time between two matching runs increases. Note that reneging cannot normally be prevented by legal means as, in most countries, it is not possible to make organ donation a legally binding obligation (Dickerson et al., 2012a, ). Both Gentry et al., (2009) and Ashlagi et al., 2011b compared KEPs with NEADs to KEPs where the final donor in a chain donates to a DDWL and obtained conflicting results. The former concluded that allowing NEAD chains decreased the overall number of transplants due to reneging risks. In contrast, the latter found that this was not the case and that, in fact, NEAD chains increased the number of transplants, including for highly sensitised recipients. The two sets of authors discussed the matter further (Ashlagi et al., 2011a, ; Gentry and Segev, , 2011), suggesting that the benefit from NEAD chains is dependent on the number of periods simulated, with a benefit shown if 8 periods are simulated (as in Ashlagi et al., 2011b ) but NEAD chains having no effect when 24 periods are simulated as in Gentry et al., (2009). This is a good time to remind the reader that many of the conclusions drawn from empirical experiments in the KEP literature are, indeed, highly dependent on the modelling assumptions used in those experiments, a fact that is usually acknowledged by the authors of such studies.
2.5 Uncertainty and robustness in KEPs
@@
Thus far, we have focused on how to select a set of exchanges for transplantation, yet in practice, some identified transplants will not proceed. In the UK’s national KEP from 2019 to 2023, around $69\%$ of selected transplants proceeded to surgery (NHS Blood and Transplant, , 2023). There are many reasons why a selected exchange may not proceed, including vertex failure (e.g., a donor or recipient becoming ill) or arc failure (e.g., a laboratory crossmatch being identified). This section gives an overview of research into uncertainty and robust optimisation in KEPs. The first direction studies stochastic approaches based on the expected likelihood of a transplant proceeding, whilst the second line of research is concerned with recourse, focusing on methods to recover transplants when parts of exchanges can no longer proceed.
@@
One proposal to account for the inevitable loss of transplants between their selection and their realisation is to consider their likelihood of failure when selecting them. A possible way of doing this is to maximise the expected number of transplants in a stochastic setting, as initially suggested by Dickerson et al., (2013) and Pedroso, (2014). The model given by Dickerson et al., (2019) provides each node and arc with a probability of success, from which a set of likely exchanges can be selected. Their branch-and-price algorithm identifies a solution that has maximum expected utility. Utilisation of such an objective can raise issues with highly-sensitised recipients not being selected, hence, Dickerson et al., (2019) extended their study to a dynamic setting that yields better outcomes for such recipients.
A different branch-and-price algorithm was provided by Alvelos et al., (2019). However, they assume that the probability of failure is equal among all nodes and equal among all arcs, which allows for internal recourse (see below in this section), while Dickerson et al., (2019) only considered the use of partially successful paths within a selected chain (i.e., when there is some failure in a chain, then the initial part of the chain can still proceed). Goldberg and Poss, (2022) extended upon this, giving a mixed-ILP model for partially successful paths within a selected chain. Whilst the models of Dickerson et al., (2019) and Alvelos et al., (2019) both require an exponential number of variables, and are solved using branch-and-price algorithms, Bidkhori et al., (2020) proposed a mixed-ILP reformulation that is compact (i.e., the numbers of variables and constraints are polynomial in the input size) and also applies when the probability of failure is inhomogeneous among arcs.
@@
In the context of stochastic matching, Blum et al., (2015) and Assadi et al., (2019) gave algorithms to approximate a maximum matching by querying the edges of a graph to determine whether they are present with some fixed probability. Along similar lines, Smeulders et al., 2022a studied the use of laboratory crossmatch tests on some potential transplants before a matching run is performed, to give certainty that those transplants may proceed. Their model thus comprises two steps. The first step selects a set of potential transplants for laboratory crossmatching. With this certainty from the laboratory tests, they then found the set of exchanges that maximised the expected number of transplants performed.
@@
Although internal recourse has been shown to improve the number of transplants that proceed in practice, other methods of recourse have been developed. For example, Klimentova et al., (2016) considered the notion of subset-recourse, where vertices for a recourse exchange can involve vertices remaining from a failed exchange as well as vertices not initially selected for transplant. Carvalho et al., (2021) studied different recourse policies and provided ILP models for each policy; their full recourse model finds a set of exchanges that maximises the number of the originally selected RDPs. Blom et al., 2024a studied recourse in terms of a three-stage defender-attacker-defender model, where the KEP organisers (the defender) select a set of exchanges, then the most disruptive set of RDPs and NDDs withdraw after being selected in a matching run (replicating an attacker or informed adversary), which is followed by the KEP repairing the solution using only the remaining RDPs and NDDs. They used a cutting plane method for the latter two stages of the problem, which allowed their observed running times to outperform the model from Carvalho et al., (2021) by an order of magnitude. Chisca et al., 2019a took a more combinatorial approach that assessed if an alternative solution could be found with minimal changes to the original selection.
@@ 2.6 Multi-Hospital and International KEPs
Klimentova et al., (2021) introduced a different credit system based on the assumption that the agents are not strategic. At a given “round” (matching run), each country has a target number of kidney transplants, representing a “fair” allocation. The difference between the actual number of transplants for a country and its target number is then used to update that country’s number of credits (positively or negatively), and the credit balances are then used to adjust the target allocations for the next round. Biró et al., (2020) conducted simulations in relation to this credit-based framework, whilst Benedek et al., 2024b extended these simulations to a larger number of countries (for cycles of length $2$ only), and considered new mechanisms for calculating credit balances. At each subsequent round, they compute a solution with the maximum number of transplants that lexicographically minimises the deviations from the target allocations for each country. The credit-based framework was extended to the case of unbounded length exchanges by Benedek et al., 2024a . Associated complexity results relating to these credit-based frameworks can be found in Biró et al., 2019b and Benedek et al., (2023, 2025).
@@
When the agents represent countries collaborating in an international KEP, additional logistical constraints may be required, such as countries permitting different exchange lengths due to differences in legislation between the countries. Mincu et al., (2021) presented an ILP model that handles these constraints in the context of an international KEP. Colley et al., (2025) extended this work to study mechanisms that are IR and IC with respect to these parameters.
@@ 2.7 Recipients’ Preferences
In a KEP, the suitability of a donor for a given recipient is generally modelled by the presence of an arc in the underlying compatibility graph and its associated cardinal weight. Indicators of the utility of a potential transplant that can contribute to this weight can include, for example, the level of HLA-matching333 Human Leukocyte Antigen matching determines the level of tissue-type compatibility between the donor and recipient. between donor and recipient, the age of the donor and the waiting time of the recipient. Instead of cardinal utilities, an alternative approach is to allow recipients to express ordinal preferences over their potential donors. In the presence of ordinal preferences, the aim is typically to find a stable set of exchanges $S$ (comprising cycles and chains), meaning that there is no blocking exchange, i.e., an exchange $E$ such that each recipient in $E$ prefers their donor in $E$ to the donor they receive in $S$ (if any).
@@
Biró and McDermid, (2010) defined the $b$-way stable $l$-way exchange problem, which is the variant of KE-Opt with ordinal recipient preferences where we seek a set of exchanges comprising cycles of length at most $l$, such that there is no blocking cycle of length at most $b$. In the case that $b=l=2$, we obtain the classical Stable Roommates problem (Irving, , 1985), as observed by Roth et al., (2005). Hence, Irving’s algorithm (Irving, , 1985) can be used to find a stable set of exchanges or report that none exists in linear time. If $b=l=\infty$, we obtain the classical Housing Market problem as observed by Roth et al., (2004), for which a stable set of exchanges always exists and can be found in linear time using Gale’s Top Trading Cycles Mechanism (Shapley and Scarf, , 1974). On the other hand, when b=l=3b=l=3Biró and McDermid, (2010) showed that the problem of deciding whether a stable set of exchanges exists is $\mathcal{NP}$-complete. Irving, (2007) showed that $\mathcal{NP}$-completeness also holds in the case that $b=3$ and $l=2$, whilst Mészáros-Karkus, (2017) proved an analogous result for the case that $b=2$ and $l=3$. Moreover, the author also showed that the problem is $\mathcal{W}[1]$-hard when parameterised by the number of cycles of length $3$.
In the same setting as Biró and McDermid, (2010), Huang, (2010) independently studied three notions of stability in KEPs that differ in their restrictiveness, namely weak, strong, and super stability. Weak stability corresponds to stability as defined informally above, and strong stability was also introduced by Biró and McDermid, (2010). Huang, (2010) showed that the $3$-way stable $3$-way exchange problem is $\mathcal{NP}$-complete for each of these notions of stability. Moreover, for so-called strong stability, Huang, (2010) showed that counting the number of sets of exchanges satisfying this property is a $\mathcal{P}-$complete problem.
Klimentova et al., (2023) gave four ILP models for the problem of finding a stable set of exchanges or reporting that none exists. They conducted simulations, measuring computation times and how many instances did not admit a stable set of exchanges. They also explored the trade-off between size (number of transplants in a set of exchanges) and stability (number of blocking exchanges). See also Biró et al., (2023); Schlotter et al., (2025); Schlotter and Mendoza-Cadena, (2025) for further results on stable sets of exchanges. Baratto et al., (2025) defined a new stability concept, which they called local stability, where a blocking exchange must contain at least one vertex from the initial set of selected exchanges.
@@
Cardinal preferences are also used in KEPs: they are most commonly represented by assigning weights to potential transplants that reflect their utility. Bertsimas et al., (2013) developed a points system that allows the KEP designer flexibility in applying fairness principles and priority criteria. A different approach was taken by Dickerson and Sandholm, (2015), who proposed a model to learn high-level objectives of a KEP from experts, and their framework implements them based on previous matching run data. Gupta et al., (2019) studied algorithmic problems connected with finding a stable set of exchanges, where stability is defined relative to cardinal valuations rather than ordinal preferences. Freedman et al., (2020) created a tie-breaking scoring function determined by a KEP’s stakeholders. Their opinions were elicited via a series of pairwise comparisons between recipient profiles, which determines what should be prioritised. Their preliminary testing of whether stakeholders’ opinions could be incorporated showed that RDPs with underdemanded blood group combinations would be negatively impacted the most. In contrast, many other RDPs’ chances would remain unchanged. The process of querying stakeholders was also followed by McElfresh et al., (2020), who proposed querying certain donors and recipients to check if they would accept a particular transplant before the selection phase. They intended to minimise rejections of selected transplants post-selection (relating to Section 2.5 on robustness in KEPs).